# HESSO: Towards Automatic, Efficient, and User Friendly Any Neural Network Training and Pruning

### v1

@@ Abstract
Structured pruning is one of the most popular approaches to effectively compress the heavy deep neural networks (DNNs) into compact sub-networks while retaining the original network performance. The existing methods suffer from multi-stage procedures along with significant engineering efforts and human expertise. The Only-Train-Once series (OTOv1-v3) has been recently proposed to resolve the many pain points by streamlining the workflow by automatically conducting (i) search space generation, (ii) structured sparse optimization, and (iii) sub-network construction. However, the built-in sparse optimizers in the OTO series, i.e., the Half-Space Projected Gradient (HSPG) family, have limitations that require hyper-parameter tuning and the implicit controls of the sparsity exploration, consequently requires intervening by human expertise. To further address such limitations, we propose a novel Hybrid Efficient Structured Sparse Optimizer (HESSO). HESSO could automatically and efficiently train a DNN within a single run to produce a high-performing sub-network. Meanwhile, it is almost tuning-free and enjoys user-friendly integration for generic training applications. To address another common issue of irreversible pruning performance collapse observed in some DNNs, we further propose a novel Corrective Redundant Identification Cycle (CRIC) to plug into HESSO for reliably identifying indispensable structures. We numerically demonstrate the efficacy of HESSO and its enhanced version HESSO-CRIC on a variety of applications ranging from computer vision to natural language processing, including large language model. The numerical results showcase that HESSO can achieve competitive performance to varying state-of-the-art benchmarks and support most DNN architectures. Meanwhile, CRIC can effectively prevent the irreversible performance collapse and further enhance the performance of HESSO on certain applications. The code is available at https://github.com/microsoft/only_train_once.
@@ 1 Introduction
Large deep neural networks (DNNs) have successfully powered a variety of applications (Ji and Chen, 2019, Zhou et al., 2024, Zhu et al., 2023). However, their typical significant time and space complexities make inference expensive and restrict deployment in resource-constrained environments. Consequently, how to compress the full DNN to the greatest extend while preserving the performance becomes essential in the many industrial and academic AI deployment pipelines. There are various model compression techniques including but not limited to pruning (Chen et al., 2021, 2023c, Fang et al., 2023), knowledge distillation (Ko et al., 2024) and quantization (Han et al., 2015), which have been well developed in the past decades.
Structured pruning typically serves as the foremost technique to produce an optimal sub-network from a pre-defined full DNN by identifying and removing redundant structures (Gale et al., 2019, Han et al., 2015, Chen et al., 2021, 2023c, Fang et al., 2023, Wang et al., 2024). Classical pruning methods focus on conducting a multi-stage procedure, requiring significant engineering efforts and expertise to manually build pruning search space, identify redundant structures, construct sub-network, and fine-tune to recover lost knowledge. To alleviate the human engineering burden, recent works (Chen et al., 2023c, b, Fang et al., 2023) have proposed pruning dependency graph to automate the pruning search space and sub-network construction. OTOv1-v2 (Chen et al., 2021, 2023c) further unify these multi-stage components together, requiring only a single training run to directly get a compact sub-network without the need of further fine-tuning. Specifically, they rely on (Dual) Half-Space Stochastic Gradient Descent (D)HSPG methods to train and prune simultaneously and have introduced a rigorous theoretical version AdaHSPG+ (Dai et al., 2023).
Figure 1: Automatic any DNN joint training and structured pruning experience achieved by the pruning mode of OTO (Chen et al., 2021, 2023c, 2023b) along with the proposed HESSO and its enhanced HESSO-CRIC optimizer. The procedure could be applied onto varying DNN and applications, and seamlessly integrated into any training pipeline to directly produce a compact pruned sub-network without need of further fine-tuning.
  DHSPG HESSO HESSO-CRIC
@@
Although OTOv1 and OTOv2 have significantly advanced the ease of use in DNN joint training and structured pruning, they still face challenges related to the complexity of the built-in (D)HSPG methods (Chen et al., 2021, 2023c, 2020c, 2020a). Specifically, these methods often require substantial hyper-parameter tuning for different downstream applications and DNN architectures. Furthermore, the sparsity explorations are implicitly controlled by (D)HSPG, which diminishes the practical convenience and usability of these approaches.
Meanwhile, many modern pruning and neural architecture search methods rely on saliency scores to identify redundant structures. However, they often suffer from performance degradation due to mistakenly identifying indispensable structures as redundant. This degradation can sometimes be irreversible due to architectural design constraints, transparency of training datasets, and the significant training resources required, posing practical challenges for their use.
To overcome these pain-points, we naturally ask, i.e., how to get a joint training and pruning optimizer which is ease-to-use, reliable, high-performing, generic and applicable onto any tasks and DNN architectures.
@@
In this work, we address this question by proposing HESSO: the Hybrid Efficient Structured Sparse Optimizer for automatic one-shot any DNN training and structured pruning. Compared to the HSPG family, HESSO offers several advantages. First, HESSO significantly simplifies the hyper-parameter setup, providing considerable practical convenience. Second, HESSO employs a progressive pruning strategy to explicitly control the sparsity exploration, making it user-friendly. Third, HESSO optionally incorporates a novel Corrective Redundancy Identification Cycle (CRIC) mechanism, so-called HESSO-CRIC, which more accurately identifies redundant groups, thereby minimizing the risk of irreversible performance collapse caused by pruning indispensable structures.
We now summarize our main contributions as follows.
@@
  • Efficient Hybrid Training and Pruning Optimizer HESSO. We propose an efficient and easy-to-use optimizer, HESSO, to enable automatic joint structured pruning and training for various model architectures and applications. HESSO progressively identifies redundant groups through flexible saliency score estimations and utilizes a hybrid training schema to effectively transfer knowledge from redundant groups to important ones, thereby maintaining the performance of the pruned model. Compared to the DHSPG in OTOv2 (Chen et al., 2023c), HESSO explicitly controls sparsity exploration, significantly reduces the need for hyper-parameter tuning, and is underpinned by theoretical foundations that ensure practical performance across various downstream applications.
@@
  • Corrective Redundancy Identification Cycle. We propose a novel Corrective Redundancy Identification Cycle (CRIC) to significantly improve the accuracy of redundancy identification. CRIC addresses the approximation errors often associated with popular Taylor-based saliency scores, thereby reducing the risk of mistakenly identifying indispensable groups as redundant. CRIC employs a voting mechanism and measures the saliency scores of each group candidate using a multi-sampling approach towards the origin. CRIC is pluggable into HESSO and helps to ensure reliable model performance by providing a more accurate assessment of group significance.
@@
  • Numerical Experiments. We validate the efficacy of HESSO across a variety of applications and model architectures. Specifically, we evaluate its performance on high-level computer vision tasks such as image classification and object detection, low-level vision tasks like super-resolution and diffusion generation, as well as natural language processing tasks including large language models. The numerical results demonstrate that HESSO performs competitively, and in many cases, exceeds the state-of-the-art benchmarks, offering significant practical convenience. Additionally, CRIC effectively mitigates the issues of irreversible collapse in pruned models, especially in challenging cases, further showcasing its utility.
@@ 2 Related Works
Automatic Structured Pruning.
Structured pruning aims to compress deep neural networks (DNNs) by removing unnecessary structures while maintaining performance (Han et al., 2015, Wen et al., 2016). The general procedure typically involves: (i) training a full model; (ii) identifying and removing redundant structures to construct a slimmer DNN based on various criteria (Lin et al., 2019, Wen et al., 2016, Li et al., 2020, Zhuang et al., 2020, Chen et al., 2017, 2018, chen2021orthant, Chen et al., 2020b, Gao et al., 2020, Zhuang et al., 2020, Meng et al., 2020, Yang et al., 2019, Zhou et al., 2019, van Baalen et al., 2020, Frankle and Carbin, 2018); and (iii) retraining the pruned model to recover any accuracy lost during pruning, with or without knowledge distillation. These methods often require a complex and time-consuming process, involving multiple training iterations and significant domain knowledge to manually handle each step.
Recent methods, such as OTO (Chen et al., 2021, 2023c, 2023b) and DepGraph (Fang et al., 2023), have made progress in automating the structured pruning process for general DNNs. OTO offers a one-shot joint training and pruning framework that can seamlessly integrate into various training processes to produce high-performing sub-networks in a single run. DepGraph, on the other hand, focuses on pruning and needs to be integrated with a multi-stage training pipeline. While these automated approaches have significantly improved user convenience, end-users still face challenges with hyper-parameter tuning and the sparse optimization expertise required to calibrate OTO’s built-in HSPG family (Chen et al., 2020c, Dai et al., 2023). Furthermore, some DNNs contain indispensable structures, the pruning of which leads to irreversible performance degradation. Identifying these critical structures remains an open problem that is often handled manually on a case-by-case basis, complicating practical use. In this work, we tackle these pain points to propose an efficient, tuning-free, and user-friendly joint training and pruning optimizer HESSO along with its enhanced version HESSO-CRIC to reliably identify indispensable structures to ensure the performance.
@@
To retain the performance of the pruned sub-network, HESSO-(CRIC) incorporates a knowledge transfer mechanism through a hybrid training schema. This approach differs from prior methods, which explicitly use knowledge distillation from unpruned models to preserve information in pruned models. Existing techniques typically require expensive computations that involve both pruned and unpruned models, whether by processing logits (Lagunas et al., 2021) or the hidden activations of intermediate layers (Xia et al., 2022, Ko et al., 2023). In contrast, our approach preserves knowledge without incurring such computational costs. Another related work, SliceGPT (Ashkboos et al., 2024), also aims to preserve computational invariance. The knowledge transfer in HESSO-CRIC similarly seeks to maintain computational invariance but does so by focusing on preserving objective function levels. However, SliceGPT is restricted to the transformer architectures and requires manually injecting additional layers. HESSO-CRIC is architecture-agnostic, efficient and user-friendly, demonstrating both scalability and versatility.
3 HESSO
Given a target DNN which variables and architecture to be optimized, HESSO formulates a constrained structured sparsity optimization problem upon the parameter groups $\mathcal{G}$ as (1).
[…]
where we seek to yield group sparsity over the prunable variables with the target sparsity level as $K$. The parameter groups can be pruning zero-invariant groups determined through the pruning dependency graph analysis or other general group formats (Chen et al., 2023c, b).
During the optimization process, HESSO begins with a warm-up phase, where the variables are trained using gradient descent or its variants, such as SGD, Adam (Kingma and Ba, 2014), or AdamW (Loshchilov and Hutter, 2017). The purpose of the warm-up stage is to collect gradient information and guide the DNN into a relatively favorable region for convergence. Following this, HESSO performs progressive pruning by periodically identifying redundant parameter groups based on predefined saliency scores. Throughout the progressive pruning phase, HESSO gradually forgets the knowledge in the redundant groups while the remaining important groups continue training, thereby facilitating the transfer and recapture of knowledge. We refer to this approach as hybrid training, where distinct training strategies are applied to different groups. Finally, once all redundant groups are identified and projected onto zero, the remaining important groups continue to be trained until final convergence. The main procedure is outlined in Algorithm 1.
@@ 3 HESSO
1:Input. Initial variable $\bm{x}_{0}$, learning rate $\alpha$, warm-up steps $T_{w}$, pruning periods $P$, period length $T_{p}$, target group sparsity level $K$, and variable partition $\mathcal{G}=\mathcal{G}_{\text{prunable}}\bigcup\mathcal{G}_{\text{unprunable}}$.
2:Warm up $T_{w}$ steps via SGD or its variants (AdamW).
3:Initialize redundant groups $\mathcal{G}_{\text{redundant}}\leftarrow\emptyset$.
4:Initialize important groups $\mathcal{G}_{\text{important}}\leftarrow\mathcal{G}$.
5:Compute sparsity level for each pruning period $\widehat{K}:=K/T_{p}$.
@@
7:     Pickup $\widehat{\mathcal{G}}_{p}$ in $\mathcal{G}_{\text{important}}$ with $\widehat{K}$-least saliency scores.
8:     Update $\mathcal{G}_{\text{redundant}}\leftarrow\mathcal{G}_{\text{redundant}}\cup\widehat{\mathcal{G}}_{p}$.
9:     Update $\mathcal{G}_{\text{important}}\leftarrow\mathcal{G}_{\text{important}}/\widehat{\mathcal{G}}_{p}$.
10:     for $t=0,1,\cdots,T_{p}-1$ do
11:         Compute trial iterate
$\widehat{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t})$ (2)
12:         Compute transferring penalty ratio for each $g\in\widehat{\mathcal{G}}_{p}$
$[\bm{\gamma}_{t}]_{g}\leftarrow\frac{T_{p}-t-1}{T_{p}-t}\frac{\left\|[\bm{x}_{t}]_{g}\right\|}{\left\|[\widehat{\bm{x}}_{t+1}]_{g}\right\|}$ (3)
13:         Update redundant groups variables
$[\bm{x}_{t+1}]_{\widehat{\mathcal{G}}_{R}}\leftarrow[\bm{\gamma}_{\widehat{\mathcal{G}}_{R}}][\widehat{\bm{x}}_{t+1}]_{\widehat{\mathcal{G}}_{R}}$ (4)
14:         Update important group variables.
$[\bm{x}_{t+1}]_{{\mathcal{G}}_{I}}\leftarrow[\widehat{\bm{x}}_{t+1}]_{{\mathcal{G}}_{I}}$ (5)
15:     end for
16:end for
17:Training important group variables till convergence.
18:Return the final iterate $\bm{x}^{*}_{\text{HESSO}}$.

### v2

# HESSO: Towards Automatic Efficient and User Friendly Any Neural Network Training and Pruning

@@ Abstract
Structured pruning is a popular technique for compressing deep neural networks (DNNs) into efficient sub-networks. However, existing methods often require multi-stage process, engineering efforts, and human expertise. The Only-Train-Once series (OTOv1-v3) has been proposed to resolve some pain points by streamlining the workflow. However, the built-in sparse optimizers in the OTO series need hyperparameter tuning and implicit control over sparsity, necessitating human intervention. To address these limitations, we propose the Hybrid Efficient Structured Sparse Optimizer (HESSO), which automatically and efficiently train a DNN within a single run to produce a high-performing sub-network. HESSO is almost tuning-free and enjoys user-friendly integration for generic training applications. In addition, to tackle the common issue of irreversible pruning performance collapse in certain DNNs, we further propose the Corrective Redundant Identification Cycle (CRIC), which integrates seamlessly with HESSO. The extensive numerical results showcase that HESSO can achieve competitive performance on various state-of-the-art benchmarks and support most DNN architectures. Moreover, CRIC can effectively prevent the irreversible performance collapse and further enhance the performance of HESSO on certain applications. The code is available at https://github.com/microsoft/only_train_once.
@@ 1 Introduction
Large deep neural networks (DNNs) have successfully powered a variety of applications (Ji and Chen, 2019, Zhou et al., 2024, Zhu et al., 2023). However, their typical significant time and space complexities make inference expensive and restrict deployment in resource-constrained environments. Consequently, how to compress the full DNN to the greatest extent while preserving the performance becomes essential in the many industrial and academic AI deployment pipelines. There are various model compression techniques including but not limited to pruning (Chen et al., 2021b, 2023c, Fang et al., 2023), knowledge distillation (Ko et al., 2024) and quantization (Han et al., 2015), which have been well developed in the past decades.
Structured pruning typically serves as the foremost technique to produce an optimal sub-network from a pre-defined full DNN by identifying and removing redundant structures (Gale et al., 2019, Han et al., 2015, Chen et al., 2021b, 2023c, Fang et al., 2023, Wang et al., 2024, Wu et al., 2024). Classical pruning methods focus on conducting a multi-stage procedure, requiring significant engineering efforts and expertise to manually build pruning search space, identify redundant structures, construct sub-network, and fine-tune to recover lost knowledge. To alleviate the human engineering burden, recent works (Chen et al., 2023c, b, Fang et al., 2023) have proposed pruning dependency graph to automate the pruning search space and sub-network construction. OTOv1-v2 (Chen et al., 2021b, 2023c) further unify these multi-stage components together, requiring only a single training run to directly get a compact sub-network without the need of further fine-tuning. Specifically, they rely on (Dual) Half-Space Stochastic Gradient Descent (D)HSPG methods to train and prune simultaneously and have introduced a rigorous theoretical version AdaHSPG+ (Dai et al., 2023).
Figure 1: Automatic any DNN joint training and structured pruning experience achieved by the pruning mode of OTO along with the proposed HESSO and its enhanced HESSO-CRIC optimizer. The procedure could be applied onto varying DNN and applications, and seamlessly integrated into any training pipeline to directly produce a compact pruned sub-network without further fine-tuning.
Although OTOv1 and OTOv2 have significantly advanced the ease of use in DNN joint training and structured pruning, they still face challenges due to the complexity of the built-in (D)HSPG methods (Chen et al., 2021b, 2023c, 2020c, 2020a). Specifically, these methods often require substantial hyper-parameter tuning for different downstream applications and DNN architectures (Dai et al., 2023, Wu et al., 2024). Furthermore, the sparsity explorations are implicit, which requires optimization expertise, thereby diminishes the practical convenience and usability.
Moreover, many modern pruning and neural architecture search methods rely on saliency scores (e.g., Taylor based) to identify redundant structures. However, they often suffer performance degradation due to mistakenly identifying indispensable structures as redundant. This degradation can sometimes be irreversible due to architectural design constraints, transparency of training datasets, and high training resource cost, posing practical challenges for their use.
  (D)HSPG HESSO HESSO-CRIC
@@
To address these issues, we propose HESSO: Hybrid Efficient Structured Sparse Optimizer for automatic one-shot any DNN training and structured pruning. Compared to the HSPG family, HESSO offers several advantages. First, it significantly simplifies the hyper-parameter setup, providing considerable practical convenience. Second, it employs a progressive pruning strategy to explicitly control the sparsity exploration, making it user-friendly. Third, HESSO optionally incorporates a novel Corrective Redundancy Identification Cycle (CRIC) mechanism, which more accurately identifies redundant groups, thereby minimizing the risk of irreversible performance collapse caused by pruning indispensable structures. We now summarize our main contributions as follows.
@@
  • Efficient Hybrid Training and Pruning Optimizer. We propose an efficient and easy-to-use optimizer, HESSO, to enable automatic joint structured pruning and training for various model architectures and applications. HESSO progressively identifies redundant groups through flexible saliency score estimations and utilizes a hybrid training schema to effectively transfer knowledge from redundant groups to important ones, thereby maintaining the performance of the pruned model. Compared to the D(HSPG) in OTO, HESSO explicitly controls sparsity exploration and knowledge transfer, minimizes the need for hyper-parameter tuning. As a result, HESSO becomes the first optimizer to realize convenient joint DNN training and pruning to the best of our knowledge.
@@
  • Corrective Redundancy Identification Cycle. We propose a novel Corrective Redundancy Identification Cycle (CRIC) to improve the accuracy of redundancy identification. CRIC addresses the approximation errors often associated with popular Taylor-based saliency scores, thereby reducing the risk of mistakenly pruning indispensable groups. CRIC employs a voting mechanism and measures the saliency scores of each group candidate using a multi-sampling approach towards the origin. CRIC can be integrated into HESSO or future joint optimizers to ensure reliable model performance by offering a more accurate assessment of group significance.
@@
  • Numerical Experiments. We validate the efficacy of HESSO and its enhanced version HESSO-CRIC across a variety of tasks. Specifically, we evaluate its performance on high-level computer vision tasks such as image classification and object detection, low-level vision tasks like super-resolution, as well as natural language processing tasks including large language models. The numerical results demonstrate that HESSO performs competitively, and in many cases, exceeds the state-of-the-art benchmarks, offering significant practical convenience. Additionally, CRIC effectively mitigates the issues of irreversible collapse in pruned models, especially in challenging cases, further showcasing its utility.
@@ 2 Related Works
In this section, we present a brief literature review on automatic structured pruning, knowledge transfer and neural architecture optimization.
General Pruning Procedures.
Structured pruning aims to compress DNNs by removing unnecessary structures while maintaining performance (Han et al., 2015, Wen et al., 2016). The general procedure typically involves: (i) training a full model; (ii) identifying and removing redundant structures to construct a slimmer DNN based on various criteria (Lin et al., 2019, He et al., 2018a, Wen et al., 2016, Li et al., 2020b, Zhuang et al., 2020, Chen et al., 2017, 2018, 2021a, 2020b, Gao et al., 2020, Zhuang et al., 2020, Meng et al., 2020, Yang et al., 2019, Zhou et al., 2019, van Baalen et al., 2020, Frankle and Carbin, 2018); and (iii) retraining the pruned model to recover any accuracy lost during pruning. These methods often require a complex and time-consuming process, involving multiple training iterations and significant domain knowledge to manually handle each step.
Automated Pruning Given Pre-defined Search Space.
To resolve the pain points of human interventions, automated pruning is raising interests from different perspectives. Given a predefined search space, AMC (He et al., 2018b) employs reinforcement learning agents to automatically determine the optimal pruning ratio. EagleEye (Li et al., 2020a) further introduces a sub-network evaluation scheme based on adaptive batch normalization, which can be integrated into AMC. OFA (Cai et al., 2020) automates the generation of sub-networks for different hardware platforms in a single process. While these approaches yield impressive performance, their application is limited to predefined search spaces. Moreover, AMC incurs additional training costs for its reinforcement learning agent. OFA’s training procedure is complex and heavy to adopt all sub-networks, which requires prior knowledge of the optimal training procedure for the largest super-network to ensure performance, making its implementation inconvenient.
Automated Pruning Over Any DNNs.
On the other hand, automatically pruning arbitrary models without prior knowledge of the search space remained a significant challenge. Recent methods, such as OTO (Chen et al., 2021b, 2023c, 2023b) and DepGraph (Fang et al., 2023), have made progress in automating the structured pruning process for general DNNs via dependency graph analysis. Subsequent works like (Wang et al., 2024) and (Ren et al., 2024) automates pruning over ONNX models. ATO (Wu et al., 2024) introduces ControlNet upon OTOv2. Among these, OTO offers a one-shot joint training and pruning framework that can seamlessly integrate into various training processes to produce high-performing sub-networks in a single run. While these automated approaches have significantly improved user convenience, end-users still face significant challenges with hyper-parameter tuning and optimization expertise to calibrate OTO’s built-in HSPG family (Chen et al., 2020c, Dai et al., 2023). Furthermore, some DNNs contain indispensable structures, the pruning of which leads to irreversible performance degradation. Identifying these critical structures remains an open problem that is often handled manually on a case-by-case basis, complicating practical use. In this work, we address these challenges by proposing an efficient, tuning-free, and user-friendly joint training and pruning optimizer, HESSO along with its enhanced version, HESSO-CRIC, which reliably identifies indispensable structures to ensure performance.
@@
To retain the performance of a pruned sub-network, HESSO-(CRIC) incorporates a knowledge transfer mechanism through a hybrid training schema. This approach differs from prior methods, which explicitly use knowledge distillation from unpruned models to preserve information in pruned models. Existing techniques typically require expensive computations that involve both pruned and unpruned models, either by processing logits (Lagunas et al., 2021) or the hidden activations of intermediate layers (Xia et al., 2022, Ko et al., 2023). In contrast, our approach preserves knowledge without incurring such computational costs. Another related works, ResRep (Ding et al., 2021b) and SliceGPT (Ashkboos et al., 2024), also aim to preserve computational invariance. The knowledge transfer in HESSO-(CRIC) similarly seeks to maintain computational invariance but does so by preserving objective function levels. However, SliceGPT is restricted to transformer architectures and requires manually injecting additional layers. ResRep is restricted to CNN architectures and require conducting structurally re-parametrization via computing resetting gradients. HESSO-(CRIC) is architecture-agnostic, efficient and user-friendly, demonstrating both scalability and versatility.
Neural Architecture Optimization.
Another related realm is the optimization over pre-specified neural architecture. NAO (Luo et al., 2018) encodes the DNN architecture into a latent representation, search over the latent space, then decodes back to a revised architecture. NAT (Guo et al., 2019) performs operator transformation upon the given DNN to produce more accurate network. These approaches transform and improve the existing DNNs, yet not search an optimal sub-network. As a result, their produced networks are typically not significantly compact compared to the baseline models. Contrarily, our approach focuses on automatically and effectively discovering compact sub-networks given pre-specified DNNs via structured pruning.
@@
3 HESSO
Given a target DNN with variables and architecture to be optimized, HESSO formulates a structured sparsity constrained optimization problem upon the set of parameter groups $\mathcal{G}$. Specifically, it aims to achieve group sparsity over the prunable variables with a target sparsity level of $K$. The optimization problem is formulated as:
[…]
where the constraint enforces that exactly $K$ parameter groups in ${\cal G}$ are pruned. These parameter groups can be defined as zero-invariant groups, which are identified through pruning dependency graph analysis, or in other general group formats (Chen et al., 2023c, b).
During the optimization process, HESSO begins with a warm-up phase, where the variables are trained using gradient descent or its variants. The purpose of the warm-up stage is to collect gradient information and guide the DNN into a relatively favorable region for convergence. Following this, HESSO performs progressive pruning by periodically identifying redundant parameter groups based on predefined saliency scores. Throughout the progressive pruning phase, HESSO gradually forgets the knowledge in the redundant groups while the remaining important groups continue training, thereby facilitating the transfer and recapture of knowledge. We refer to this approach as hybrid training, where distinct training strategies are applied to different groups. Finally, once all redundant groups are identified and projected onto zero, the remaining important groups continue to be trained until final convergence. The main procedure is outlined in Algorithm 1.
@@ 3 HESSO
1:Input. Initial variable $\bm{x}_{0}$, learning rate $\alpha$, warm-up steps $T_{w}$, pruning periods $P$, period length $T_{p}$, target group sparsity level $K$, and variable partition $\mathcal{G}=\mathcal{G}_{I}\bigcup\mathcal{G}_{R}$.
2:Warm up $T_{w}$ steps via SGD or its variants.
3:Initialize redundant groups $\mathcal{G}_{R}\leftarrow\emptyset$.
4:Initialize important groups $\mathcal{G}_{I}\leftarrow\mathcal{G}$.
5:Compute sparsity for each pruning period $\widehat{K}:=K/T_{p}$.
@@
7:     Pickup $\widehat{\mathcal{G}}_{p}$ in $\mathcal{G}_{I}$ with $\widehat{K}$-least saliency scores.
8:     Update $\mathcal{G}_{R}\leftarrow\mathcal{G}_{R}\cup\widehat{\mathcal{G}}_{p}$ and $\mathcal{G}_{I}\leftarrow\mathcal{G}_{I}/\widehat{\mathcal{G}}_{p}$.
9:     for $t=0,1,\cdots,T_{p}-1$ do
10:         Compute trial iterate $\widehat{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t}).$
11:         Compute transferring ratio for each $g\in\widehat{\mathcal{G}}_{p}$:
$[\bm{\gamma}_{t}]_{g}\leftarrow\frac{T_{p}-t-1}{T_{p}-t}\frac{\left\|[\bm{x}_{t}]_{g}\right\|}{\left\|[\widehat{\bm{x}}_{t+1}]_{g}\right\|}.$
12:         Update redundant and important variables:
$\begin{split}[\bm{x}_{t+1}]_{\widehat{\mathcal{G}}_{p}}&\leftarrow[\bm{\gamma}_{t}]_{\widehat{\mathcal{G}}_{p}}[\widehat{\bm{x}}_{t+1}]_{\widehat{\mathcal{G}}_{p}},\\ [\bm{x}_{t+1}]_{{\mathcal{G}}_{I}}&\leftarrow[\widehat{\bm{x}}_{t+1}]_{{\mathcal{G}}_{I}}.\end{split}$
13:     end for
14:end for
15:Training important group variables till convergence.
16:Return the final iterate $\bm{x}^{*}_{\text{HESSO}}$.